The current position of the harmonic oscillator is given by the relation 

x(t)=(0,3m)cos(πt/2).Time t is expressed in seconds. Find the period of the oscillator.



1
Expert's answer
2021-03-04T11:53:52-0500

"\omega=\dfrac{\pi}{2}rad/s^{-1}"

"T=\dfrac{2\pi}{\omega}"

"\Rightarrow T=\dfrac{2\pi\times2}{\pi}\\\Rightarrow T=4\space s"